Demand D = 18,500 units/year


Ordering cost S = $26/order


Holding cost H = $4/unit/year



a) Calculate the EOQ for the workbooks.


b) What are the annual holding costs for the workbooks?


c) What are the annual ordering costs?



1
Expert's answer
2021-11-11T03:24:02-0500

Daily demand = 60000/240 = 250 If lead time = 3 days (lead time < time between orders) Reorder point = (60000/240)x3=750 Reorder when inventory on hand = 750